OpenStudy (anonymous):

Using Lagrange Multipier, maximize/minimize f(x,y,z)=cos(x)cos(y)cos(z) subject to x+y+z=pi

OpenStudy (anonymous):

f = cos(x)cos(y)cos(z) let g= x+y+z - pi ( from the constraint ) now we require the grad f = scalar multiple of g let the scalar be Y ie grad f = Y grad g

OpenStudy (anonymous):

grad f = ( -sin(x)cos(y)cos(z) , -cos(x)sin(y)cos(z) , -cos(x)cos(y)sin(z) ) grad g = ( 1,1,1)

OpenStudy (anonymous):

therefore -sin(x)cos(y)cos(z) = Y -cos(x)sin(y)cos(z) = Y -cos(x)cos(y)sin(z) = Y